OpenStudy (anonymous):

\int\limits_0^pi/2 cos4x dx

OpenStudy (y1066380):

\[\int\limits\limits_0^\pi2 \cos4x dx \] ?

OpenStudy (y1066380):

The answer is 0

OpenStudy (anonymous):

can u show me the process?:)

OpenStudy (y1066380):

\[2\int\limits\limits_0^\pi \cos4x dx \]

OpenStudy (y1066380):

wait, it is difficult to type it out

OpenStudy (anonymous):

okay :)

OpenStudy (y1066380):